Two cars start from two towns 700 miles apart and travel towards each other. They pass each other after 7 hours. Find the speed of each car if the difference between their speeds is 60 miles per hour.
Answer by checkley79(3341)

D=RT
LET X BE THE SPEED OF THE FASTER CAR
LET X+60 BE THE SPEED OF THE FASTER CAR.
700=(X+X+60)7
700=(2X+60)7
700=14X+420
14X=700-420
14X=280
X=280/14
X=20 MPH FOR THE SLOWER CAR.
20+60=80 MPH FOR THE FASTER CAR.
PROOF:
700=(20+80)7
700=100*7
700=700